in the sequence 1,22,4444,88888888,....then 1025th term will be?

Grade:12

3 Answers

V. Harshit -
8 years ago
i dont think the 1025 term will fit in thebox
Akshita Goyal
8 years ago
@v harshit no the answer is 2^10.
 
Akshay
8 years ago
@Akshita : Yes you are write 1025th term will be 2^10.
sequence= {1,2,2,4,4,4,4,8,8,8,8,8,8,8,8......}
Note that terms from 2i  to term number (2i+1-1) have the value of 2^i.
Ex: i=2, 2^2=4 and 2^3-1=7. So, term number 4 to 7 are = 4
//ly i=3, 2^3=8 and 2^4-1=15. so, term number 8 to 15 are =8.
Similarly for i=4, term number 16 to 31 are =16.
i=10, term number 1024 to 2047 will be equal = 2^10.
Sequence is like {1 repeated 1 times, 2^1 repeated 2^1 times, 2^2 repeated 2^2 times, 2^3 repeated 2^3 times and so on}.
I believe you have done a mistake in putting commas(,)  in the question